# Towards Better Benchmark Datasets for Inductive Knowledge Graph Completion

### v2

@@ Abstract
Knowledge Graph Completion (KGC) attempts to predict missing facts in a Knowledge Graph (KG). Recently, there’s been an increased focus on designing KGC methods that can excel in the inductive setting, where a portion or all of the entities and relations seen in inference are unobserved during training. Numerous benchmark datasets have been proposed for inductive KGC, all of which are subsets of existing KGs used for transductive KGC. However, we find that the current procedure for constructing inductive KGC datasets inadvertently creates a shortcut that can be exploited even while disregarding the relational information. Specifically, we observe that the Personalized PageRank (PPR) score can achieve strong or near SOTA performance on most inductive datasets. In this paper, we study the root cause of this problem. Using these insights, we propose an alternative strategy for constructing inductive KGC datasets that helps mitigate the PPR shortcut. We then benchmark multiple popular methods using the newly constructed datasets and analyze their performance. The new benchmark datasets help promote a better understanding of the capabilities and challenges of inductive KGC by removing any shortcuts that obfuscate performance.
1 Introduction
@@
Knowledge Graph Completion (KGC) attempts to predict unseen facts given an existing knowledge graph (KG). KGC has numerous applications including drug discovery [1], personalized medicine [2], and recommendations [3]. Traditionally, most research on KGC was focused on the transductive setting, where the same sets of entities and relations are seen during training and testing. Most methods [4, 5, 6] generally focus on learning embeddings for all entities and relations to facilitate the prediction of new facts.
@@
In recent years, interest in KGC has shifted towards designing methods that can generalize to new entities or relations not seen during training. This task, known as “inductive KGC”, requires a method to train on a graph $\mathcal{G}_{\text{train}}$ and perform inference on a different graph $\mathcal{G}_{\text{inf}}$, where the inference graph contains either new entities, relations, or both. Because of this, methods for inductive KGC don’t rely on fixed embeddings for entities or relations, instead opting for more flexible techniques that can inductively learn representations based on a given graph [7, 8, 9]. To asses the ability of methods for this task, new datasets have been constructed that require methods to reason inductively. All inductive datasets [7, 10, 9] are constructed from existing transductive KGC datasets. This is done by sampling two graphs, one each for train and inference, which contain disjoint entities. Multiple methods [8, 11, 9] have reported tremendous promise on these newer benchmark datasets.
@@
However, we find that on almost all inductive datasets, we can achieve competitive performance by using the Personalized PageRank [12] (PPR) score to perform inference. We note that PPR is a non-learnable heuristic and ignores the relational information in the graph. In Figure 1, we compare the performance of PPR against the supervised SOTA performance on both inductive and transductive datasets. We can see that when performing KGC on inference graphs with either new entities (E) or new entities and relations (E, R), PPR performs only roughly 25% worse than SOTA. However, this is generally not true for transductive datasets, where PPR usually performs poorly. Interestingly, this observation is true for inductive datasets even when the transductive dataset that it is created from has poor PPR performance. For example, while PPR performs very poorly on FB15k-237, it achieves higher performance on it’s inductive derivatives (denoted by “FB”). These findings are problematic as PPR has no basis in literature as a heuristic for KGC, since it completely overlooks the relational aspect of KGs. Therefore, this suggests the potential existence of a shortcut that allows a simple non-learnable method like PPR to achieve high performance on almost all inductive datasets. This also brings into question how successful most methods are in inductive reasoning, as a large portion of their performance may be due to this shortcut.
@@
Figure 1: Hits@10 of PPR vs. Supervised SOTA. Results are on (a) (E) Inductive, (b) (E, R) Inductive, and (c) Transductive datasets.
2 Background and Related Work
@@
All existing inductive datasets are sampled from existing transductive datasets. Furthermore, the majority of inductive datasets [7, 9] are further created via the same procedure introduced by [7]. We now give a brief overview of this procedure. Given a transductive dataset, which we denote as $\mathcal{G}$, $k$ seed entities are randomly chosen from $\mathcal{G}$. The 2-hop neighborhood is then extracted around each individual seed entity. To prevent exponential growth, the number of neighbors sampled at any hop is capped at 50 for each seed entity. The resulting edges are then combined to create $\mathcal{G}_{\text{train}}$ and are subsequently removed from the original graph. The inference graph is then sampled in a similar manner from the resulting graph $\mathcal{G}\setminus\mathcal{G}_{\text{train}}$. One exception to this procedure are the ILPC datasets, introduced by [10], which instead sample $p\%$ of the nodes from $\mathcal{G}$ and use them to create $\mathcal{G}_{\text{train}}$. The rest of the nodes are then used to construct $\mathcal{G}_{\text{inf}}$. In practice, we find that both methods tend to produce similar graphs.
@@
Inductive KGC Methods: NeurlLP [13] and DRUM [14] consider combining the path representations of different length between both entities in a triple. However, since they explicitly consider each path, they are often limited to only considering paths of up to length 2 or 3. Conditional MPNNs [15] are a more efficient alternative to encoding higher-order path information. They work by conditioning the message passing mechanism on the known entity, allowing the implicit aggregation of all paths up to a length $L$ (which is equal to the number of GNN layers). The value of $L$ is typically $5/6$. Prominent examples include NBFNet [8] and RED-GNN [11]. More scalable alternatives have been proposed that prune the propagated messages [16, 17, 18]. NodePiece [19] is concerned with parameter efficiency, using an anchor-based approach to learn a more compact set of entity and relation representations. All previous methods assume that a fixed set of relations exist between the train and inference graphs. To account for new relations in inference, InGRAM [9] introduces the concept of a “relation graph”, which inductively encodes the representation of each relation. Note that we omit subgraph methods that have been to shown to prohibitively expensive [7, 20, 21, 22, 23] or those that require the use of external textual information [24, 25].
@@
Personalized PageRank: PageRank [12] computes the probability of finishing a random walk of arbitrary length at some node $u$, when there is equal probability of beginning the walk at any node in the graph. Personalized PageRank (PPR) [12] is a version of PageRank that is “personalized" to some root node $s$, where at each step there is a probability $\alpha$ of teleporting to $s$. The set of PPR scores for a root node $s$ is given by $\text{pr}_{s}\in\mathbb{R}^{\lvert\mathcal{V}\rvert}$ and can be formulated as the weighted sum of all random walk probabilities between two nodes [26]:
$\text{pr}_{s}=\alpha\sum_{k=0}^{\infty}(1-\alpha)^{k}W^{k}x_{s},$ (1)
@@
where $W$ is the random walk matrix and $x_{s}$ is a one-hot vector at node $s$. Observe that the weight given to a walk decays with the increase in length due to $(1-\alpha)^{k}$. As such, the PPR score will often be higher for those nodes of shorter distance to $s$. For a KG, we obtain the PPR matrix by first converting the inference graph, $\mathcal{G}_{\text{inf}}$, to an undirected graph. This is common practice in KGC [27] whereby inverse relations are added to the graph. We further assume an edge weight of 1 for all edges. As such, the relations are completely disregarded when computing the PPR.
3 Preliminary Study
@@
3.1 Personalized Pagerank (PPR) is a Strong Baseline for Inductive KGC
@@
(b) $\Delta\text{SPD}$ vs. PPR Hits@10
Figure 2: (a) Percent increase in PPR performance on the inductive datasets as compared to their parent transductive dataset. For datasets with multiple splits (i.e., FB and WN) we take the mean performance across each. (b) Relationship between $\Delta\text{SPD}$ and performance. We observe a strong relationship among both transductive and inductive datasets.
@@
In Figure 1, we show the performance when using the PPR versus the SOTA performance among supervised methods. The SOTA method is dataset dependent, with the specific methods listed in Appendix A.1. We split the datasets by transductive, (E), and (E, R) inductive. We include those datasets most often used in each task, comprising 25 in total. Please see Appendix B.1 for more on the datasets chosen. The full set of results can also be found in Appendix A.1.
We observe that on both types of inductive datasets, the PPR score does reasonably well, performing on average only 25-29% less than SOTA. This is surprising as PPR is both non-learnable and ignores the relational aspect of the KG. We also find that on some datasets like the WN or ILPC inductive splits, the performance nearly matches or exceeds the supervised performance. On the other hand, for the transductive datasets, the performance disparity is often much larger. Interestingly, we note that PPR still performs well on some transductive datasets, including the popular WN18RR. This tells us that this phenomenon is not necessarily unique to inductive datasets, but is most apparent there.
@@
3.2 When Does PPR Perform Well and Why?
@@
We find that the answer lies in comparing the mean shortest path distance (SPD) between positive and negative samples. For a query $(s,r,*)$, let’s denote $o^{+}$ as the correct answer and $o^{-}\in\mathcal{V}^{-}(s,r)$ as the set of negative answers for the query. We compute the SPD on the inference graph for both the $o^{+}$ and $\mathcal{V}^{-}(s,r)$. This is repeated for each test sample $(s,q,*)\in\mathcal{E}_{\text{inf}}$. We then calculate the mean SPD across all positive and negative test samples, which we denote as $\text{SPD}^{+}$ and $\text{SPD}^{-}$, respectively. The different in mean SPD is correspondingly given by $\Delta\text{SPD}=\text{SPD}^{+}-\text{SPD}^{-}$. This tells us, on average, how much shorter the distance between entities in positive samples are relative to those in negatives.
In Figure 2(b) we show the relationship between $\Delta\text{SPD}$ and the Hits@10 when using PPR. The results are across 25 different datasets, for both inductive and transductive KGC. Calculating the Pearson correlation, we can observe it is 0.87, which indicates a strong relationship between the two metrics. Therefore, there exists a basic pattern to distinguish positive and negative samples in many KGC datasets. Put simply, the SPD between entities in positive test samples tend to be lower than that in negative samples. The larger the discrepancy between the mean distances, the better PPR can perform. This suggests that the PPR scores can exploit this pattern in the datasets, to achieve good performance, even while completely ignoring the relational information.
@@
Note that we use PPR as opposed to SPD in our experiments since all-pairs SPD is costly to calculate and the full PPR matrix can be efficiently approximated via [28]. Furthermore, as we show in Section 3.3, even when controlling for the SPD, the PPR score can help differentiate between positive and negative samples.
3.3 Why is This So Common on Inductive Datasets?
(a) FB15k-237
(b) WN18RR
Figure 3: The mean shortest path (SP) distance between positive test samples and negative samples. We display the results for two transductive datasets (FB15k-237 and WN18RR) and the inductive datasets that are derived from them.
Observation 1: The current procedure for creating inductive datasets increases the $\Delta\text{SPD}$ and thereby the performance of PPR. As noted earlier, all common inductive datasets are created from existing transductive datasets. See Section 2 for a detailed overview of the construction process. In a nutshell, the training and inference graph are constructed sequentially by sampling a number of subgraphs from a graph. In Figure 3 we show the mean SPD of both positive and negative samples for 12 inductive datasets and their parent transductive dataset. We limit our analysis to those datasets derived from FB15k-237 and WN18RR, as the majority of inductive datasets are derived from them. We observe that for all inductive datasets, while the mean SPD for negative samples sharply rises, the mean SPD for positive samples stay roughly the same. This creates the shortcut described in Section 3.2, where the SPD can easily differentiate between positive and negative samples. Since the gap between the distances is almost always larger than the original transductive dataset, this shortcut becomes more pronounced, thereby leading to a better PPR performance (see Appendix A.1 for detailed results).
Table 1: % Increase in Positive vs. Negative sample PPR, broken down by SPD.

### v3

@@ Abstract.
Knowledge Graph Completion (KGC) attempts to predict missing facts in a Knowledge Graph (KG). Recently, there’s been an increased focus on designing KGC methods that can excel in the inductive setting, where a portion or all of the entities and relations seen in inference are unobserved during training. Numerous benchmark datasets have been proposed for inductive KGC, all of which are subsets of existing KGs used for transductive KGC. However, we find that the current procedure for constructing inductive KGC datasets inadvertently creates a shortcut that can be exploited even while disregarding the relational information. Specifically, we observe that the Personalized PageRank (PPR) score can achieve strong or near SOTA performance on most datasets. In this paper, we study the root cause of this problem. Using these insights, we propose an alternative strategy for constructing inductive KGC datasets that helps mitigate the PPR shortcut. We then benchmark multiple popular methods using the newly constructed datasets and analyze their performance. The new benchmark datasets help promote a better understanding of the capabilities and challenges of inductive KGC by removing any shortcuts that obfuscate performance. The code and dataset and can be found https://github.com/HarryShomer/Better-Inductive-KGC.
Knowledge graphs, Link prediction
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737438isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Machine learning
1. Introduction
@@
Knowledge Graph Completion (KGC) attempts to predict unseen facts given an existing knowledge graph (KG). KGC has numerous applications including drug discovery (Zeng et al., 2022), personalized medicine (Chandak et al., 2023), and recommendations (Wang et al., 2021). Traditionally, most research on KGC was focused on the transductive setting, where the same sets of entities and relations are seen during training and testing. Most methods (Bordes et al., 2013; Trouillon et al., 2016; Schlichtkrull et al., 2018) generally focus on learning embeddings for all entities and relations to facilitate the prediction of new facts.
@@
In recent years, interest in KGC has shifted towards designing methods that can generalize to new entities or relations not seen during training. This task, known as “inductive KGC”, requires a method to train on a graph $\mathcal{G}_{\text{train}}$ and perform inference on a different graph $\mathcal{G}_{\text{inf}}$, where the inference graph contains either new entities, relations, or both. Because of this, methods for inductive KGC don’t rely on fixed embeddings for entities or relations, instead opting for more flexible techniques that can inductively learn representations based on a given graph (Teru et al., 2020; Zhu et al., 2021; Lee et al., 2023). To assess the ability of methods for this task, new datasets have been constructed that require methods to reason inductively. All inductive datasets (Teru et al., 2020; Galkin et al., 2022; Lee et al., 2023) are constructed from existing transductive KGC datasets. This is done by sampling two graphs, one each for train and inference, which contain disjoint entities. Multiple methods (Zhu et al., 2021; Zhang and Yao, 2022; Lee et al., 2023) have reported tremendous promise on these newer benchmark datasets.
@@
However, we find that on almost all inductive datasets, we can achieve competitive performance by using the Personalized PageRank (PPR) score (Page et al., 1999) to perform inference. We note that PPR is a non-learnable heuristic and ignores the relational information in the graph. In Figure 1, we compare the performance of PPR against the supervised SOTA performance on both inductive and transductive datasets. We can see that when performing KGC on inference graphs with either new entities (E) or new entities and relations (E, R), PPR performs only roughly 25% worse than SOTA. However, this is generally not true for transductive datasets, where PPR usually performs poorly. Interestingly, this observation is true for inductive datasets even when the transductive dataset that it is created from has poor PPR performance. For example, while PPR performs very poorly on FB15k-237, it achieves higher performance on it’s inductive derivatives (denoted by “FB”). These findings are problematic as PPR has no basis in literature as a heuristic for KGC, since it completely overlooks the relational aspect of KGs. Therefore, this suggests the potential existence of a shortcut that allows a simple non-learnable method like PPR to achieve high performance on almost all inductive datasets. This also brings into question how successful most methods are in inductive reasoning, as a large portion of their performance may be due to this shortcut.
@@
Figure 1. Hits@10 of PPR vs. Supervised SOTA. Results are on (a) (E) Inductive, (b) (E, R) Inductive, and (c) Transductive datasets.
2. Background and Related Work
@@
All existing inductive datasets are sampled from existing transductive datasets. Furthermore, the majority of inductive datasets (Teru et al., 2020; Lee et al., 2023) are further created via the same procedure introduced by (Teru et al., 2020). We now give a brief overview of this procedure. Given a transductive dataset, which we denote as $\mathcal{G}$, $k$ seed entities are randomly chosen from $\mathcal{G}$. The 2-hop neighborhood is then extracted around each individual seed entity. To prevent exponential growth, the number of neighbors sampled at any hop is capped at 50 for each seed entity. The resulting edges are then combined to create $\mathcal{G}_{\text{train}}$ and are subsequently removed from the original graph. The inference graph is then sampled in a similar manner from the resulting graph $\mathcal{G}\setminus\mathcal{G}_{\text{train}}$. One exception to this procedure are the ILPC datasets, introduced by (Galkin et al., 2022), which instead sample $p\%$ of the nodes from $\mathcal{G}$ and use them to create $\mathcal{G}_{\text{train}}$. The rest of the nodes are then used to construct $\mathcal{G}_{\text{inf}}$. In practice, we find that both methods tend to produce similar graphs.
@@
Inductive KGC Methods: NeurlLP (Yang et al., 2017) and DRUM (Sadeghian et al., 2019) consider combining the path representations of different length between both entities in a triple. However, since they explicitly consider each path, they are often limited to only considering paths of up to length 2 or 3. Conditional MPNNs (Huang et al., 2024) are a more efficient alternative to encoding higher-order path information. They work by conditioning the message passing mechanism on the known entity, allowing the implicit aggregation of all paths up to a length $L$ (which is equal to the number of GNN layers). The value of $L$ is typically $5$ or $6$. Prominent examples include NBFNet (Zhu et al., 2021) and RED-GNN (Zhang and Yao, 2022). More scalable alternatives have been proposed that prune the propagated messages (Zhu et al., 2023; Zhang et al., 2023; Shomer et al., 2023b). NodePiece (Galkin et al., 2021) is concerned with parameter efficiency, using an anchor-based approach to learn a more compact set of entity and relation representations. All previous methods assume that a fixed set of relations exist between the train and inference graphs. To account for new relations in inference, InGRAM (Lee et al., 2023) introduces the concept of a “relation graph”, which inductively encodes the representation of each relation. Gao et al. (2023) introduces the concept of “double permutation- equivariant representations” as a way to model KGs that are equivariant to permutations of both the entities and relations. They theoretically show that capturing this property is essential for proper generalization across KGs. To this point these introduce a new method, ISDEA/ISDEA+, that can satisfy this property. They further introduce a variant of InGRAM (Lee et al., 2023), DEq-InGram, that endows it with the ability to compute double equivariant representations. Note that we omit subgraph methods that have been to shown to prohibitively expensive (Teru et al., 2020; Liu et al., 2021; Mai et al., 2021; Xu et al., 2022; Geng et al., 2023) or those that require the use of external textual information (Gesese et al., 2022; Daza et al., 2021).
@@
Personalized PageRank: PageRank (Page et al., 1999) computes the probability of finishing a random walk of arbitrary length at some node $u$, when there is equal probability of beginning the walk at any node in the graph. Personalized PageRank (PPR) (Page et al., 1999) is a version of PageRank that is “personalized" to some root node $s$, where at each step there is a probability $\alpha$ of teleporting to $s$. The set of PPR scores for a root node $s$ is given by $\text{pr}_{s}\in\mathbb{R}^{\lvert\mathcal{V}\rvert}$ and can be formulated as the weighted sum of all random walk probabilities between two nodes (Chung, 2007):
(1) $\text{pr}_{s}=\alpha\sum_{k=0}^{\infty}(1-\alpha)^{k}W^{k}x_{s},$
@@
where $W=D^{-1}A$ and $x_{s}$ is a one-hot vector at node $s$. Observe that the weight given to a walk decays with the increase in length due to $(1-\alpha)^{k}$. As such, the PPR score will often be higher for those nodes of shorter distance to $s$. For a KG, we obtain the PPR matrix by first converting the inference graph, $\mathcal{G}_{\text{inf}}$, to an undirected graph. This is common practice in KGC (Dettmers et al., 2018) whereby inverse relations are added to the graph. We further assume an edge weight of 1 for all edges. As such, the relations are completely disregarded when computing the PPR.
3. Preliminary Study
@@
3.1. Personalized Pagerank (PPR) is a Strong Baseline for Inductive KGC
@@
(b) $\Delta\text{SPD}$ vs. PPR Hits@10 for all datasets
Figure 2. (a) Percent increase in PPR performance on the inductive datasets as compared to their parent transductive dataset. For datasets with multiple splits (i.e., FB and WN) we take the mean performance across each. (b) Relationship between $\Delta\text{SPD}$ and performance. We observe a strong relationship among both transductive and inductive datasets.
@@
In Figure 1, we show the performance when using the PPR versus the SOTA performance among supervised methods. The SOTA method is dataset dependent, with the specific methods are shown in Appendix A.1.. We split the datasets by transductive, (E), and (E, R) inductive. We include those datasets most often used in each task, comprising 25 in total. Please see Appendix B.1 for more on the datasets chosen. The full set of results are shown in Table 6.
We observe that on both types of inductive datasets, the PPR score does reasonably well, performing on average only 25-29% less than SOTA. This is surprising as PPR is both non-learnable and ignores the relations of the KG. We also find that on some datasets like the WN or ILPC inductive splits, the performance nearly matches or exceeds the supervised performance. On the other hand, for the transductive datasets, the performance disparity is often much larger. Interestingly, we note that PPR still performs well on some transductive datasets, including the popular WN18RR. This tells us that this phenomenon is not necessarily unique to inductive datasets, but is most apparent there.
@@
3.2. When Does PPR Perform Well and Why?
@@
We find that the answer lies in comparing the mean shortest path distance (SPD) between positive and negative samples. For a query $(s,r,*)$, let’s denote $o^{+}$ as the correct answer and $o^{-}\in\mathcal{V}^{-}(s,r)$ as the set of negative answers for the query. We compute the SPD on the inference graph for both the $o^{+}$ and $\mathcal{V}^{-}(s,r)$. This is repeated for each test sample $(s,q,*)\in\mathcal{E}_{\text{inf}}$. We then calculate the mean SPD across all positive and negative test samples, which we denote as $\text{SPD}^{+}$ and $\text{SPD}^{-}$, respectively. The difference in mean SPD is correspondingly given by $\Delta\text{SPD}=\text{SPD}^{-}-\text{SPD}^{+}$. This tells us, on average, how much shorter the distance between entities in positive samples are relative to those in negatives.
In Figure 2(b) we show the relationship between $\Delta\text{SPD}$ and the Hits@10 when using PPR. The results are across 25 different datasets, for both inductive and transductive KGC. Calculating the Pearson correlation, we can observe it is 0.87, which indicates a strong relationship between the two metrics. Therefore, there exists a basic pattern to distinguish positive and negative samples in many KGC datasets. Put simply, the SPD between entities in positive test samples tend to be lower than that in negative samples. The larger the discrepancy between the mean distances, the better PPR can perform. This suggests that the PPR scores can exploit this pattern in the datasets, to achieve good performance, even while completely ignoring the relation information.
@@
Table 1. % Increase in Positive vs. Negative sample PPR, broken down by SPD.
@@
Note that we use PPR as opposed to SPD in our experiments since all-pairs SPD is costly to calculate and the full PPR matrix can be efficiently approximated via (Andersen et al., 2006). Furthermore, as we show in Section 3.3, even when controlling for the SPD, the PPR score can help differentiate between positive and negative samples.
3.3. Why is This So Common on Inductive Datasets?
(a) FB15k-237
(b) WN18RR
Figure 3. The mean shortest path (SP) distance between positive test samples and negative samples. We display the results for two transductive datasets (FB15k-237 and WN18RR) and the inductive datasets that are derived from them.
Observation 1: The current procedure for creating inductive datasets increases the $\Delta\text{SPD}$ and thereby the performance of PPR. As noted earlier, all common inductive datasets are created from existing transductive datasets. See Section 2 for a detailed overview of the construction process. In a nutshell, the training and inference graph are constructed sequentially by sampling a number of subgraphs from a graph. In Figure 3 we show the mean SPD of both positive and negative samples for 12 inductive datasets and their parent transductive dataset. We limit our analysis to those datasets derived from FB15k-237 and WN18RR, as the majority of inductive datasets are derived from them. We observe that for all inductive datasets, while the mean SPD for negative samples sharply rises, the mean SPD for positive samples stay roughly the same. This creates the shortcut described in Section 3.2, where the SPD can easily differentiate between positive and negative samples. Since the gap between the distances is almost always larger than the original transductive dataset, this shortcut becomes more pronounced, thereby leading to a better PPR performance.